# SReferences

### v1

@@ III-A Backbone
Following PolarNet [24], we first convert the raw point cloud $\mathbf{P}\in\mathbb{R}^{N\times K}$ into a polar bird’s-eye view (BEV) grid of size $H\times W$, where $H$ and $W$ correspond to the radial and angular dimensions, respectively. Each grid cell accumulates points based on their 3D polar coordinates. We then employ a simplified PointNet-style module [49], which applies a series of linear transformations interleaved with batch normalization and ReLU activation. After processing each grid cell, we aggregate features along the vertical $Z$-dimension using max pooling, producing a fixed-size representation $\mathbf{M}\in\mathbb{R}^{H\times W\times C}$, where $C=512$ is the feature dimension. We adopt polar rather than Cartesian coordinates to ensure a more balanced distribution of points over different ranges, helping the network focus on closer, denser areas without excessive quantization loss [24]. To capture multi‐scale context, we employ a U‐Net‐style encoder [50] that progressively downsamples the BEV representation by a factor of 2×2$\times$2 ×, 4×4$\times$4 ×, 8×8$\times$8 ×, and 16×16$\times$16 ×. These feature maps are fed to the decoders, allowing the network to integrate both coarse global context and fine local details in subsequent segmentation and instance‐level tasks.
@@ IV-B Training Protocol
[…]
@@ IV-C Benchmarking Results
In Tab. II, we further assess our method on nuScenes [21]. Our method leverages training unknown labels to guide uncertainty modeling, which outperforms methods based on K+1 supervision alone. Similarly, in Tab. III, we report performances on KITTI-360 using our defined Vocabulary Unseen. We observe a similar trend, with our model consistently outperforming all baselines for both known and unknown panoptic segmentation. These results underscore the advantages of integrating uncertainty modeling with structured feature learning. Our uncertainty-driven losses enable the network to effectively delineate the boundaries between known and unknown regions, leading to a more robust and balanced segmentation performance across both known and unknown objects.
@@ IV-D2 Impact of Proposed Losses
Tab. V presents the impact of our proposed uncertainty losses on the LiDAR open-set panoptic segmentation performance across two vocabulary settings. We first compare two baseline mechanisms for handling unknowns. The Proto-Unknowns baseline, following [17], is implemented in our ULOPS by removing the uncertainty estimation component and introducing a learnable global constant $U$, which serves as the score for not associating with any known prototype. In contrast, the ULOPS implicitly identifies unknowns through uncertainty estimation. Notably, the Association-Based Unknowns baseline yields a PQ of 58.4% and 57.1% with UQ at 15.1% and 23.2% on Vocabulary Unseen and Vocabulary 1 respectively. Meanwhile, the ULOPS model demonstrates an improved known-unknwon separation (UQ of 27.9% and 31.2%), while also achieving improvement in PQ, indicating that leveraging uncertainty cues can generalize to unseen classes even without explicit unknown annotations. However, exposure to unknown samples significantly improves this separation. Incorporating Uniform Evidence Loss increases UQ (+3.8% on Vocabulary Unseen, +3.3% on Vocabulary 1), showing that enforcing a flat Dirichlet distribution for unknown voxels effectively raises per-voxel uncertainty. However, PQ remains relatively unchanged, suggesting that while the network becomes more uncertain about unknowns, it does not necessarily refine segmentation quality. Following, introducing Adaptive Uncertainty Separation Loss further enhances UQ (+0.6% and +0.6% over Uniform Evidence Loss), confirming that leveraging batch-level statistics ensures an improved distinction between known and unknown uncertainties. Notably, this improvement comes without a significant PQ drop, suggesting that the loss dynamically balances uncertainty separation without overly penalizing confident predictions for known objects. Finally, Contrastive Uncertainty Loss achieves the most substantial gains, boosting UQ to 35.3% (+7.4%) and 38.6% (+7.4%) across the two vocabularies while also improving PQ. This validates our hypothesis that explicitly enforcing a margin-based separation between known and unknown uncertainties leads to a more structured and meaningful uncertainty landscape. The simultaneous PQ improvement suggests that this contrastive approach also enhances feature representations.

### v2

@@ III-A Backbone
Following PolarNet [24], we first convert the raw point cloud $\mathbf{P}\in\mathbb{R}^{N\times K}$ into a polar bird’s-eye view (BEV) grid of size $H\times W$, where $H$ and $W$ correspond to the radial and angular dimensions, respectively. Each grid cell accumulates points based on their 3D polar coordinates. We then employ a simplified PointNet-style module [49], which applies a series of linear transformations interleaved with batch normalization and ReLU activation. After processing each grid cell, we aggregate features along the vertical $Z$-dimension using max pooling, producing a fixed-size representation $\mathbf{M}\in\mathbb{R}^{H\times W\times C}$, where $C=512$ is the feature dimension. We adopt polar rather than Cartesian coordinates to ensure a more balanced distribution of points over different ranges, helping the network focus on closer, denser areas without excessive quantization loss [24]. To capture multi‐scale context, we employ a U‐Net‐style encoder [50] that progressively downsamples the BEV representation by a factor of $2\times$, $4\times$, $8\times$, and $16\times$. These feature maps are fed to the decoders, allowing the network to integrate both coarse global context and fine local details in subsequent segmentation and instance‐level tasks.
@@ IV-B Training Protocol
We discretize the 3D space to $[480,360,32]$ voxels for SemanticKITTI ($r:1\sim$60\text{\,}\mathrm{m}$,z:-3\sim$3\text{\,}\mathrm{m}$\$) and for nuScenes ($r:0\sim$50\text{\,}\mathrm{m}$,z:-5\sim$3\text{\,}\mathrm{m}$\$), where $r$ denotes the radial distance from the LiDAR sensor. The network is trained for 60 epochs with a batch size of 4 and a crop size of $[240,180,32]$. We use the Adam optimizer with an initial learning rate of $0.01$, with stepwise decay by a factor of 10 at 45 and 55 epochs. We apply instance augmentation along with random point cloud augmentations (flipping, rotation, scaling). We set $t=3$ for the uncertainty threshold and $F=32$ for the embedding size. Loss weights are set as follows: $\mathcal{L}_{\text{center}}=200$, $\mathcal{L}_{\text{uniform}}=0.1$, $\mathcal{L}_{\text{adaptive}}=0.1$, and $\mathcal{L}_{\text{contrastive}}=0.7$, with all other losses weighted at 1.
@@ IV-C Benchmarking Results
In Tab. II, we further assess our method on nuScenes [21]. Our method leverages training unknown labels to guide uncertainty modeling, which outperforms methods based on K+1 supervision alone. Similarly, in Tab. X, we report performances on KITTI-360 using our defined Vocabulary Unseen. We observe a similar trend, with our model consistently outperforming all baselines for both known and unknown panoptic segmentation. These results underscore the advantages of integrating uncertainty modeling with structured feature learning. Our uncertainty-driven losses enable the network to effectively delineate the boundaries between known and unknown regions, leading to a more robust and balanced segmentation performance across both known and unknown objects.
@@ IV-D2 Impact of Proposed Losses
Tab. VIII presents the impact of our proposed uncertainty losses on the LiDAR open-set panoptic segmentation performance across two vocabulary settings. We first compare two baseline mechanisms for handling unknowns. The Proto-Unknowns baseline, following [17], is implemented in our ULOPS by removing the uncertainty estimation component and introducing a learnable global constant $U$, which serves as the score for not associating with any known prototype. In contrast, the ULOPS implicitly identifies unknowns through uncertainty estimation. Notably, the Association-Based Unknowns baseline yields a PQ of 58.4% and 57.1% with UQ at 15.1% and 23.2% on Vocabulary Unseen and Vocabulary 1 respectively. Meanwhile, the ULOPS model demonstrates an improved known-unknwon separation (UQ of 27.9% and 31.2%), while also achieving improvement in PQ, indicating that leveraging uncertainty cues can generalize to unseen classes even without explicit unknown annotations. However, exposure to unknown samples significantly improves this separation. Incorporating Uniform Evidence Loss increases UQ (+3.8% on Vocabulary Unseen, +3.3% on Vocabulary 1), showing that enforcing a flat Dirichlet distribution for unknown voxels effectively raises per-voxel uncertainty. However, PQ remains relatively unchanged, suggesting that while the network becomes more uncertain about unknowns, it does not necessarily refine segmentation quality. Following, introducing Adaptive Uncertainty Separation Loss further enhances UQ (+0.6% and +0.6% over Uniform Evidence Loss), confirming that leveraging batch-level statistics ensures an improved distinction between known and unknown uncertainties. Notably, this improvement comes without a significant PQ drop, suggesting that the loss dynamically balances uncertainty separation without overly penalizing confident predictions for known objects. Finally, Contrastive Uncertainty Loss achieves the most substantial gains, boosting UQ to 35.3% (+7.4%) and 38.6% (+7.4%) across the two vocabularies while also improving PQ. This validates our hypothesis that explicitly enforcing a margin-based separation between known and unknown uncertainties leads to a more structured and meaningful uncertainty landscape. The simultaneous PQ improvement suggests that this contrastive approach also enhances feature representations.
@@ IV-E Qualitative Results
Table VI: Imapact of uncertainity-driven losses on KITTI-360 performance.
Method mIoU (t+1) mIoU (t)
OccFormer + COTR - 41.2
Baseline 33.9 38.2
PeakOcc 37.3 41.6
Table VII: Imapact of uncertainity-driven losses on KITTI-360 performance.
Method mIoU (t+1) mIoU (t)
OccFormer - 12.1
Baseline 6.0 7.2
PeakOcc 8.7 11.8
$u=1-\exp\left(-\frac{1}{N}\sum_{i=1}^{N}\|f-f_{\text{aug}_{i}}\|_{2}\right)$ (10)
$f_{\text{fused}}=w_{\text{cam}}\cdot f_{\text{cam}}+w_{\text{lidar}}\cdot f_{\text{lidar}}$ (11)
Table VIII: Imapact of uncertainity-driven losses on KITTI-360 performance.
Method PQ UQ PQ UQ
Proto-Unknowns [17] 58.4 15.1 57.1 23.2
ULOPS 60.4 27.9 59.1 31.2
+ Uniform Evidence 61.0 31.7 59.7 34.5
+ Adaptive Uncertainty 61.1 32.3 59.9 35.1
+ Contrastive Uncertainty 61.5 35.3 60.6 38.6
Table IX: Performance comparison on KITTI-360 with our proposed Vocabulary Unseen as described in Sec. IV-A.
Method Modality PQ PQTh PQSt
Efficient-LPS L 62.0 56.8 70.6
SMAC-Seg HiRes L 68.4 68.0 68.8
CPSeg HR L 71.1 71.5 70.6
Ours L 72.3 72.3 73.5
Panoptic-PHNet LC 74.7 74.0 75.9
Ours LC 76.4 77.2 75.3
Table X: Performance comparison on KITTI-360 with our proposed Vocabulary Unseen as described in Sec. IV-A.
Number of Augmentations PQ PQTh PQSt
- 74.8 75.1 74.5
1 75.4 75.8 74.8
3 76.4 77.2 75.3
5 75.9 76.4 75.1